NEW RESULTS FROM THE ROTATIONAL SPECTRA OF THE HALOGEN MONOXIDES. INTERATOMIC POTENTIALS, FINE AND HYPERFINE INTERACTIONS.

\underline{B. J. DROUIN}, E. A. COHEN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109-8099; C. E. MILLER, Department of Chemistry, Haverford College, Haverford, PA 19041-1392; and H. S. P. MÜLLER, I. Physikalisches Institut, Universität zu Köln, Zülpicher Str. 77, D-50937, Köln, Germany.

Last year we reported the IO rotational spectrum in vibrational levels up to v = 13 for the X_{1} ^{2}Π_{3/2} state and to v = 9 for the X_{2} ^{2}Π_{1/2} state. In comparison, previous studies of BrO were limited to v = 2 for X_{1} and v = 1 for the X_{2} state. Using the DC discharge cell that was so effective for vibrational excitation of IO, measurements of BrO have now been extended up to v = 8 and v = 7 for the X_{1} and X_{2} states, respectively. Excited vibrational levels of ClO were not obtained with these methods, however good signal to noise at thermal populations allowed measurements for the v = 2 levels of both the X_{1} and X_{2} states as well as ^{18}O in natural abundance. The Hamiltonian of Brown, et. al.^{a}, with explicit isotope dependencies for each parameter, has provided a set of mass and nuclear moment independent parameters for each of the halogen monoxide species. The electron spin-rotation constant, γ, and the centrifugal distortion of the spin-orbit splitting, A_{D}, which are normally correlated, have been separately determined by the isotope dependence of their contributions to the spectrum. Interatomic potentials have been derived from the mass-independent parameters that are accurate up to the observed excitation energies for each molecule. Analyses of the fine-structure parameters indicate that these molecules are close to the single perturber limit. The hyperfine parameters will be compared with the literature values^{b}^{,}^{c}^{,}^{d} of the appropriate calculated relativistic radial integrals of the halogens.

aJ. M. Brown and J. K. G. Watson, J. Mol. Spec. 148, 371 (1991).

bP. Pyykkö, M. Seth, Theor. Chem. Acc. 96, 92-104 (1997).

cI. Lindgren and A. Rosén, Case Studies in Atomic Physics, 4, 197-298 (1974).

dP. Pyykkö and L. Wiesenfeld, Mol. Phys. 43, 557-580 (1981).